Mir Status 20 October 1997

Date: Monday, October 20, 1997 // MD 76/25

Mir Station System/Subsystem Activities

The second IVA into Spektr was conducted today, and the crew connected two of the three cables for the solar array orientation system to the hermaplate on the hatch. They were unable to make one of the connections due to difficulties with access, so one of the three undamaged arrays will not be steerable. The MOST is attempting to find out which one this is, but it is believed to be the array opposite the damaged array. The current plan is to make the connection to the Kristall avionics on Wednesday, the 22nd.

Mr. Blagov told us that they do not yet know the cause of the failure to depressurize the Progress vestibule on Saturday. They suspect that a limit switch was not making contact, preventing the supply of power to the valve. However, the depressurization worked fine on Sunday. He said that the crew will take a closer look for causes when they reopen Progress.

Mir Payload Activities

SPEKTR IVA
MG:FAC:I:39 BEM/BSTC-M & BTR Facility Maintenance

IVA Powered Hardware Operations

Network Status - A total loss of communication, including telephone, fax, e-mail, and voice loop occurred at 1300 DMT today. Marshall opened two trouble tickets to determine the anomaly. Communication between the TsUP and the US was restored at approximately 20:00 DMT, and the anomaly was determined to be a loss of both circuits due to a faulty memory card on Node 214.

Experiment Hardware

BSTC-M - Dave reported that the temperatures of the first two chambers were 34 and 35 degrees, and the setpoint was still at 36 degrees. He also mentioned that the air inlet temperature is warmer and up to 32.4 degrees since the BSTC is in a warmer spot now (has been around 29).

Environmental Parameters (previous 24 hour period):

Parameter Min. Max. Nominal Range
P (mmHg) BB 723 728 660 - 850
ppO2 (mmHg) 165 168 140 - 200
ppCO2 (mmHg) 3.7 4.1 up to 8.0
ppH20 (mmHg) 13.5 13.9 8 - 14
Humidity (%) 47 52 30 - 75
Temp (xC) BB 27.3 28.4 18 - 28

Communications Summary

A/G with Consultant Group - No dedicated comm today due to IVA.

ISSUES - Due to the IVA running slightly overtime, there will be no telemetry available this evening. This will prohibit the re-powering of the Priroda module and therefore the CAPE/MIM equipment will not be re-powered as planned. This will be the first activity tomorrow post Priroda power-up.

During the 20:59-21:14 Wallops comm pass the link from Mir to ground was not successful. Wallops could hear the crew, but due to problems in Moscow the MCC-M could not hear the crew. During the next pass, comm was nominal.

  • submit to reddit


Loading



Wernher von Braun Memorial Symposium September 10 - 12, 2019
Global MilSatCom, November 5-7, 2019, London, UK




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on October 20, 1997 4:30 PM.

Mir Status 19 October 1997 was the previous entry in this blog.

Mir Status Report 19 November 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.